Find the area of a triangle if b=5 ft., h=2 ft.
Answer by misscrt(37)

Find the area of a triangle if b=5 ft., h=2 ft.
Formula for area of triangle is 1/2bh = A
So, plug the values for b and h into the formula to get:
1/2(5)(2) = A
10/2 = A
5 =A
Remember the units for area,A, is ft^2
A = 5 ft^2